Can you show me the work and answer for 2x-3.5=x-3.5 and finding the value of x making the equation true. Thank you
Answer by ankor@dixie-net.com(22740)

2x-3.5=x-3.5
Find x
2x - 3.5 = x - 3.5
subtract x from both sides
2x - x - 3.5 = -3.5
add 3.5 to both sides
x = 0
Replace x with 0 in the equation and you will see equality reigns!